loganengle1423:

Michelle rents a movie for a flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Michelle has the movie.

Isolation:

So, we need a y=mx+b function, where m is slope/rate, and b is the y-intercept where it starts we see, "flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie", this means that 1.5 is the y-intercept, b, and 1.25, the rate, is slope get the equation?

Hoodmemes:

I'v found the answer options

1 attachment
jbug2004:

tiff blocked me lol

Hoodmemes:

(I'm giving a baby explanation so deal with it) Ok so as we can see it says "Michelle rents a movie for a flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie." So she already has to rent the movie for $1.50 right? "$1.50 plus an additional $1.25" This is a hint that all we have to do is add 1.50 to 1.25 (we ain't done yet tho) \[1.50+1.25 \] And Additional Money for (HOW MANY NIGHTS SHE RENTS THIS MOVIE) So it depends on how many nights she keeps this movie rental to decide how many times we will (Multiply) (1.25) (An x would come after the 1.25) We can now choose what answer options we can cross off. We can cross off (1.50x + 1.25) due to the fact that in the worded problem she already states she rented the movie for (1.50) dollars it does not depend on how many days she keeps this movie for that certain fee, so we wouldn't even have an x-(Will not be Multiplying) 1.50 at all . We can also cross off (2.75) bc we can't just add the two number mentioned in the problem together,the problem states we have to do way more than that (It basically was stating that 1.25 would be multiplied depending on how many nights she keeps the movie rental)-"I'v stated this before"- so we couldn't just add the two number together and call that an answer. So (1.50x + 1.25) and (2.75) would be incorrect . This leaves answer options ( (1.50 + 1.25)x ) and (1.50 + 1.25x) (Remember as i said we are only Multiplying (1.25 due to the fact that "it will be multiplied depending on how long she keeps the movie rental) With that being said we can basically cross off (1.50 + 1.25)x bc that means we are multiplying both 1.50 plus 1.25 by how long she keeps the movie rental (It ain't even said we had to do all that in the problem) SO your answer would most likely be { c (x)= (1.50 + 1.25x ) } This isn't my best work.

Isolation:

that's not hte answer

Hoodmemes:

Elaborate more bc by typing { c (x)= (1.50 + 1.25x ) } I basically meant x= 1.50 + 1.25x

Isolation:

oh sorryyyyyyyyyyyyyy, with all those parenthesis and stuff, I thought you put the parentheses before teh x, meaning (1.5+1.25)x, sorry~

Hoodmemes:

its ight I make common mistakes as such too